# External Prompt Features Enhanced Parameter-efficient Fine-tuning for Salient Object Detection

### v1

@@ Abstract
Salient object detection (SOD) aims at finding the most salient objects in images and outputs pixel-level binary masks. Transformer-based methods achieve promising performance due to their global semantic understanding, crucial for identifying salient objects. However, these models tend to be large and require numerous training parameters. To better harness the potential of transformers for SOD, we propose a novel parameter-efficient fine-tuning method aimed at reducing the number of training parameters while enhancing the salient object detection capability. Our model, termed EXternal Prompt features Enhanced adapteR Tuning (ExPert), features an encoder-decoder structure with adapters and injectors interspersed between the layers of a frozen transformer encoder. The adapter modules adapt the pre-trained backbone to SOD while the injector modules incorporate external prompt features to enhance the awareness of salient objects. Comprehensive experiments demonstrate the superiority of our method. Surpassing former state-of-the-art (SOTA) models across five SOD datasets, ExPert achieves 0.215 mean absolute error (MAE) in ECSSD dataset with 80.2M trained parameters, 21% better than transformer-based SOTA model and 47% better than CNN-based SOTA model.
@@ 1.1 Motivation
Salient object detection (SOD) is a widely studied task in computer vision that outputs a binary mask of the visual salient objects in an image. The detection of salient objects can benefit various other computer vision tasks, such as semantic segmentation, instance segmentation, and object detection. In recent years, convolutional neural network (CNN) based models and transformer-based models have shown promising performance for SOD. However, although transformer-based models [26, 20, 14] generally outperform their CNN counterparts, they are more computationally expensive due to their typically large number of parameters that are essential for achieving superior performance.
The encoder-decoder framework is widely used for salient object detection, which is defined as a binary semantic segmentation. Firstly, a vision encoder is initialized with pre-trained model weights from classification or segmentation models. The next step is to fine-tune the encoder and decoder on salient object detection datasets to extend the model to SOD task. The predicted salient masks are generated by the specific decoder with the extracted features. Beyond fine-tuning, training pre-trained backbone along with other new sophisticated modules can gain better performances. However, it necessitates an even larger number of trained parameters.
@@ 1.2 Methods Overview
We propose EXternal Prompt features Enhanced adapteR Tuning (ExPert) model to parameter-efficiently tune pre-trained transfomer backbones for salient object detection. ExPert is a backbone-agnostic model and can be extended to any transformer-based pre-trained backbones. Inspired by [4, 16], ExPert uses the block-level [^1] In [4], the adapter is a side connection of the feed forward function inside transformer block which is denoted as ”FFT-level”. In [16] and our ExPert, the adapter is a side connection between transformer block and is denoted as ”block-level”. adapter module to tune the transformer backbone between each block unit. We denote the adapter of ExPert as E-adapter.
@@
Moreover, we hypothesize that the captions of images are highly related to the salient elements. Based on this premise, ExPert interacts BLIP’s visual features and text embeddings of corresponding captions using cross attention. The best result was achieved by injecting the interacted features combined with ViT’s features to the backbone. Comprehensive experiments show that ExPert surpasses CNN-based SOTA models largely and performs better than previous transformer-based models.
@@ 1.3 Contributions
    We propose ExPert model to parameter-efficiently fine-tune pre-trained transformer backbones for salient object detection. ExPert is backbone-agnostic and can use different transformer-based backbones. Comprehensive experiments demonstrate the superiority of ExPert.
@@
    We design the block-level E-adapter to parameter-efficiently adapt the pre-trained transformer backbones to salient object detection. The size of trained parameters of ExPert is only 80.2M.
@@
    Our E-injector can receive different external prompt features and inject them to guide the backbone to extract salient features. Experiments demonstrate that the injection of features that contain rich semantic information boosts the performance largely.
@@ 2.1 Salient Object Detection
CNN-based models are proficient at extracting local details. EGNet [27] focuses on the complementarity between edge and content of salient objects by extracting edge information. U2Net [17] proposes a nested U-shape convolutional network to handle input with flexible sizes without any pre-training. Although requiring more computing costs, transformer-based models surpass CNN-based models in SOD because transformer can grasp the long-range semantic context of input images. SelfReformer [26] adopts a global branch to refine the local context branch with a multi-stage transformer backbone to achieve better long-range information extraction. EVP [16] fine-tunes SegFormer [22] with patch embedding prompt and Fourier transformation prompt to better differentiate objects.
@@ 2.2 Adapter Tuning and Visual Prompt Tuning
Adapter tuning is a method to fine-tune pre-trained models which was firstly proposed in [19] as a trainable side connection branch for parameter-efficient tuning. Later Houlsby et al. [9] used adapter tuning to parameter-efficiently train transformer-based language models. AdapterFormer [4] applies adapter tuning to Vision Transformer and achieves promising performance in multi-label classification. EVP [16] demonstrates that adapter tuning can effectively transfer pre-trained transformer-based models to downstream tasks such as salient object detection, camouflaged object detection, and other binary segmentation tasks.
Prompt [15] is originally used in natural language processing (NLP) to instruct pre-trained language model to understand and shift to new tasks. Prompt tuning has also developed rapidly in computer vision (CV) domain. An input-agnostic visual perturbation prompt is learned and fed to a model together with input images to repurpose pre-trained models to downstream tasks in [2]. Some learnable parameters are injected into transformer’s input space to efficiently fine-tune large-scale transformer models in [10]. ViT-Adapter [5] uses side branches to inject spatial prior into ViT to fine-tune the model for detection and segmentation tasks. These works all show that the injection of prompt information to the original backbone can guide pre-trained models for versatile downstream tasks.
@@ 3.1 Overview
Salient object detection is an important task in computer vision field, which detects the most salient objects in an RGB image and outputs binary masks of these objects. Let $I\in\mathbb{R}^{H\times W\times 3}$ denote the input image and $G\in\mathbb{R}^{H\times W\times 1}$ the corresponding ground truth binary mask. The output binary mask of model is $M\in\mathbb{R}^{H\times W\times 1}$. Suppose the SOD model is $\mathcal{F}$ and its parameters as $\theta$, then the mask is calculated as $\mathcal{F}(I,\theta)$. The loss function $\mathcal{L}$ in ExPert is a combination of binary cross entropy (BCE) loss and intersection over union (IoU) loss [25]. The training target is to minimize $\mathcal{L}(M,G)$ between $M$ and $G$.
We propose an encoder-decoder model denominated as EXternal Prompt features Enhanced adapteR Tuning (ExPert) with block-level E-adapter and E-injector. The architecture of ExPert is shown in Fig. 1 and entails a vision encoder, a mask decoder, some E-adapter modules between the transformer blocks of vision encoder and some shared E-injectors for each feature scale.
Since SOD is defined as segmentation task which is similar to semantic segmentation, pre-trained transformer backbones for segmentation or classification are preferable. A multi-scale encoder of SegFormer [22] and its decoder are chosen as the backbone and the segmentation head of ExPert. E-adapter is a lightweight side connection module that helps to transfer pre-trained transformer backbone to salient object detection. In addition, E-injector is a lightweight side connection that projects vision features from other backbones as guiding prompts and injects these prompts into the encoder. The detailed structures of E-adapter and E-injector are in Fig. 3.
ExPert is trained in an end-to-end manner with image-mask pairs. During training, the vision encoder is frozen while the E-adapters, E-injectors and decoder are trained. Experiments demonstrate that the combined prompts of BLIP and ViT achieve the best result. As a backbone-agnostic model, ExPert can switch between different transformer backbones with simple modification while the decoder need to be specified according to different backbones.
@@ 3.2 Encoder and Decoder
Figure 2: The decoder of ExPert for multi-scale features. The illustration feature images is visualized by choosing a random slice of channel dimension. ExPert’s final mask is generated by resizing this mask to the original size.
Transformer backbones can be classified into two types according to whether the scale of features changes. One is the single-scale backbone and the other is the multi-scale backbone. The features of single-scale backbone keep the same size during the forward propagation while the multi-scale backbone’s features change size. For salient object detection, former researches [28, 17, 26] emphasized the importance of multi-scale feature fusion to get finer segmentation mask. Since multi-scale features is crucial to dense prediction for finer details, multi-scale backbones have an advantage over single-scale backbones in segmentation tasks. Therefore we choose the the MiT-B4 version of SegFormer [22] as our multi-scale backbone. The encoder of SegFormer has 4 stages of feature extraction and each stage has different scales of features. The decoder is kept intact as the original one as shown in Fig. 2.
For each stage of feature extraction, denote the output feature of $i_{th}$ stage as $F^{out}_{i}$. A linear layer $L_{i}$ projects $F^{out}_{i}$ to align the dimension followed by a bilinear upsampling $\mathcal{U}()$ to align the scale. Then all aligned features from different stages are concatenated by $Cat()$ to get $F_{c}$. Finally, two convolution layers, $C_{fuse}$, $C_{pred}$ and sigmoid function $\mathcal{S}()$ are applied for the mask generation. The mask $M$ is generated as Eq. 2.
@@ 3.3 E-adapter
Figure 3: The detailed structure of E-adapter and E-injector. D-P is down projection, M-L is median linear layer, U-P is up projection and Dim-P is dimension projection.
To fine-tune the transformer backbone in a parameter-efficient manner, we employ the adapter tuning method, which reduces the feature dimensionality through a bottleneck design, thereby diminishing the number of trained parameters. As shown in the bottom branch of Fig. 3, we propose the E-adapter, consisting of a down-projection layer with parameters $P^{d}_{e}$, a median linear layer $L^{m}_{e}$ and a up-projection layer with parameters $P^{u}_{e}$. It is noteworthy that the down-projection of E-adapter is shared for features of the same scale in order to further diminish the number of trained parameters. The $L^{m}_{e}$ and up-projection is independent for each E-adapter. The additional low-dimension $L^{m}_{e}$ serves to increase the variability of E-adapter. Define the $i_{th}$ block forward function as $\mathcal{B}_{i}()$, the $(i+1)_{th}$ block’s feature $F_{i+1}$ after the E-adapter is computed as Eq. 4.
@@ 3.4 E-injector
Previous works [5, 16] show that side connection module like adapter can introduce extra information to boost the model’s performance in object detection and segmentation tasks. We design the E-injector to inject external features from other backbones of same input images as guiding prompts into the encoder for salient object detection. E-injector’s structure is depicted in the top branch of Fig. 3.
@@
If the number of injected prompt features is $J$, the $j_{th}$ E-injector is composed of a prompt transformation $\mathcal{T}_{j}$ and a dimension projection layer $P^{dim}_{j}$. Since visual prompts might vary in size and shape, a feature transformation $\mathcal{T}_{j}$ fits the prompt feature $F^{\prime}_{j}$ to $i_{th}$ layer’s feature $F_{i}$ of the frozen backbone. $\mathcal{T}_{j}$ is composed of a normalization and a resize operation. E-injector can receive different transformer features $F^{\prime}_{j}$ including the features from DINO, ViT and BLIP’s vision encoder. The output of E-injector $F^{inj}_{j}$ is generated as Eq. 5. To better adjust the injector features to the backbone, a learnable scaling vector $\alpha_{j}$ is used to weight them. The $(i+1)_{th}$ block’s feature $F_{i+1}$ is computed as Eq. 6.
$F^{inj}_{j}$ $=P^{dim}_{j}\cdot\mathcal{T}_{j}(F^{\prime}_{j})$ (5)
$F_{i+1}$ $=\mathcal{B}_{i}(F_{i})+F^{ad}_{i}+\sum_{j=1}^{J}F^{inj}_{j}\times\alpha_{j}$ (6)
@@
Finding suitable prompt features is crucial to the salient masks quality for E-injector. DINO [3] is a self-supervised model trained without labels that exhibits an obvious tendency to focus on objects in an image. Observing the attention maps of DINO reveals coarse masks that are nearly similar to the masks of the salient object, we assume this kind of object-aware features can aide pre-trained models to locate objects. Besides, considering that our multi-scale backbone has fewer layers of the full-size scale compared to ViT, our backbone’s whole perception of an image might be complemented by the features of ViT’s last layer. ExPert uses the features of ViT/B-16 as the auxiliary features for better global perception of images.
@@
While piling up the features might strengthen the visual details, it is still hard to guide the model to recognize the notion of saliency. It is noteworthy that the caption naturally contains the descriptions of the salient objects in an image which are beneficial to SOD. Therefore, we envision that the caption of an image is highly related to the salient objects and contains the salient information. Some Vision Language Models (VLM) like BLIP [12] are trained on large caption datasets. [^2] CLIP [18] is a well-know VLM model trained with millions of image-text pairs, however the text of CLIP is a simple sentence with the class name which is not the caption of the whole image. And the resolution of CLIP’s training images is 224*224, the feature is 7*7 with patch size of 32 which is too small to upsample. Therefore ExPert does not consider CLIP’s features. The features from BLIP’s vision encoder trained with image-text labels, which contains rich semantic information of an image, is injected as semantic-enhanced features by E-injector.
@@
To fully explore the rich semantic information in BLIP, we interact BLIP’s vision features and BLIP’s captions for better focus of salient objects. For an image $I$, the corresponding caption is generated by BLIP using beam search. The caption is tokenized and embedded by BLIP’s text encoder to get the text embedding $T$. The interacted feature is acquired via the cross-attention between the last layer vision feature $V_{b}$ of BLIP and $T$. In each cross attention layer, $V_{b}$ is projected as query and $T$ is projected as key and value by linear projections. Experiments show that one cross-attention layer is enough for the interaction, more details can be found in the supplement file.
@@
Since the features from ViT and similar model can refine the mask in details and the features from BLIP can inject semantic information to ExPert, we combine these two features together for E-injector. The final version of ExPert uses the feature from ViT’s last layer and the interacted feature by cross-attention from BLIP. Experiments show that the combination prompt version achieves the best performance. Unless specified, ExPert represents our best version.

### v2

@@ Abstract
Salient object detection (SOD) aims at finding the most salient objects in images and outputs pixel-level binary masks. Transformer-based methods achieve promising performance due to their global semantic understanding, crucial for identifying salient objects. However, these models tend to be large and require numerous training parameters. To better harness the potential of transformers for SOD, we propose a novel parameter-efficient fine-tuning method aimed at reducing the number of training parameters while enhancing the salient object detection capability. Our model, termed EXternal Prompt features Enhanced adapteR Tuning (ExPert), features an encoder-decoder structure with adapters and injectors interspersed between the layers of a frozen transformer encoder. The adapter modules adapt the pre-trained backbone to SOD while the injector modules incorporate external prompt features to enhance the awareness of salient objects. Comprehensive experiments demonstrate the superiority of our method. Surpassing former state-of-the-art (SOTA) models across five SOD datasets, ExPert achieves 0.215 mean absolute error (MAE) in the ECSSD dataset with 80.2M trained parameters, 21% better than SelfReformer[32] and 47% better than EGNet[34].
@@ 1.1 Motivation
Salient object detection (SOD) is a widely studied task in computer vision that outputs a binary mask of the visually salient objects in an image. The detection of salient objects can benefit various computer vision tasks, such as semantic segmentation, instance segmentation, and object detection. In recent years, convolutional neural network (CNN) based models and transformer-based models have shown promising performances for SOD. However, although transformer-based models [32, 22, 15] generally outperform their CNN counterparts, they are more computationally expensive due to their typically large number of parameters that are essential for achieving superior performance.
The encoder-decoder framework is widely used for salient object detection, which is defined as a binary semantic segmentation. Firstly, a vision encoder is initialized with pre-trained model weights from classification or segmentation models. The next step is to fine-tune the encoder and decoder on salient object detection datasets to extend the model to the SOD task. The predicted salient masks are generated by the specific decoder with the extracted features. Beyond fine-tuning, training the pre-trained backbone along with other new sophisticated modules can gain better performance. However, it necessitates an even larger number of trained parameters.
@@ 1.2 Methods Overview
We propose EXternal Prompt features Enhanced adapteR Tuning (ExPert) model to parameter-efficiently tune pre-trained transformer backbones for salient object detection. ExPert is a backbone-agnostic model and can be extended to any transformer-based pre-trained backbones. Inspired by [4, 17], ExPert uses the block-level [^1] In [4], the adapter is a side connection of the feed-forward function inside the transformer block which is denoted as ”FFT-level”. In [17] and our ExPert, the adapter is a side connection between transformer blocks and is denoted as ”block-level”. adapter module to tune the transformer backbone between each block unit. We denote the adapter of ExPert as E-adapter.
@@
Moreover, we hypothesize that the captions of images are highly related to the salient elements. Based on this premise, ExPert interacts BLIP’s visual features and text embeddings of corresponding captions using cross attention. The best result was achieved by injecting the interacted features combined with ViT’s features into the backbone. Comprehensive experiments show that ExPert surpasses CNN-based SOTA models largely and performs better than previous transformer-based models.
@@ 1.3 Contributions
    We propose the ExPert model to parameter-efficiently fine-tune pre-trained transformer backbones for salient object detection. ExPert is backbone-agnostic and can use different transformer-based backbones. Comprehensive experiments demonstrate the superiority of ExPert.
@@
    We design the block-level E-adapter to parameter-efficiently adapt the pre-trained transformer backbones to salient object detection. The size of the trained parameters of ExPert is only 80.2M.
@@
    Our E-injector can receive different external prompt features and inject them to guide the backbone to extract salient features. Experiments demonstrate that the injection of features that contain rich semantic information largely boosts the performance.
@@ 2.1 Salient Object Detection
CNN-based models are proficient at extracting local details. EGNet [34] focuses on the complementarity between edges and the content of salient objects by extracting edge information. U2Net [19] proposes a nested U-shape convolutional network to handle inputs with flexible sizes without any pre-training. Although requiring more computing costs, transformer-based models surpass CNN-based models in SOD because transformer models can grasp the long-range semantic context of input images. SelfReformer [32] adopts a global branch to refine the local context branch with a multi-stage transformer backbone to achieve better long-range information extraction. EVP [17] fine-tunes SegFormer [27] with patch embedding prompts and Fourier transformation prompts to better differentiate objects.
@@ 2.2 Adapter Tuning and Visual Prompt Tuning
Adapter tuning is a method to fine-tune pre-trained models which was first proposed in [21] as a trainable side connection branch for parameter-efficient tuning. Later Houlsby et al. [9] used adapter tuning to parameter-efficiently train transformer-based language models. AdapterFormer [4] applies adapter tuning to Vision Transformer and achieves promising performance in multi-label classification. EVP [17] demonstrates that adapter tuning can effectively transfer pre-trained transformer-based models to downstream tasks such as salient object detection, camouflaged object detection, and other binary segmentation tasks.
Prompt [16] is originally used in natural language processing (NLP) to instruct pre-trained language models to understand and shift to new tasks. Prompt tuning has also developed rapidly in the computer vision (CV) domain. An input-agnostic visual perturbation prompt is learned and fed to a model together with input images to repurpose pre-trained models to downstream tasks in [2]. Some learnable parameters are injected into the transformer’s input space to efficiently fine-tune large-scale transformer models in [10]. ViT-Adapter [5] uses side branches to inject spatial priors into ViT to fine-tune the model for detection and segmentation tasks. These works all show that the injection of prompt information into the original backbone can guide pre-trained models for versatile downstream tasks.
@@ 3.1 Overview
Salient object detection is an important task in the computer vision field, which detects the most salient objects in an RGB image and outputs binary masks of these objects. Let $I\in\mathbb{R}^{H\times W\times 3}$ denote the input image and $G\in\mathbb{R}^{H\times W\times 1}$ the corresponding ground truth binary mask. The output binary mask of the model is $M\in\mathbb{R}^{H\times W\times 1}$. Suppose the SOD model is $\mathcal{F}$ and its parameters are $\theta$, then the mask is calculated as $\mathcal{F}(I,\theta)$. The loss function $\mathcal{L}$ in ExPert is a combination of binary cross entropy (BCE) loss and intersection over union (IoU) loss [30]. The training target is to minimize $\mathcal{L}(M,G)$ between $M$ and $G$.
We propose an encoder-decoder model denominated as EXternal Prompt features Enhanced adapteR Tuning (ExPert) with block-level E-adapter and E-injector. The architecture of ExPert is shown in Fig. 1 and entails a vision encoder, a mask decoder, some E-adapter modules between the transformer blocks of the vision encoder and some shared E-injectors for each feature scale.
Since SOD is defined as a segmentation task that is similar to semantic segmentation, pre-trained transformer backbones for segmentation or classification are preferable. A multi-scale encoder of SegFormer [27] and its decoder are chosen as the backbone and the segmentation head of ExPert. E-adapter is a lightweight side connection module that helps to transfer the pre-trained transformer backbone to salient object detection. In addition, E-injector is a lightweight side connection module that projects vision features from other backbones as guiding prompts and injects these prompts into the encoder. The detailed structures of E-adapter and E-injector are illustrated in Fig. 3.
ExPert is trained in an end-to-end manner with image-mask pairs. During training, the vision encoder is frozen while the E-adapters, E-injectors and decoder are trained. Experiments demonstrate that the combined prompts of BLIP and ViT achieve the best result. As a backbone-agnostic model, ExPert can switch between different transformer backbones with simple modification [^2] More details can be found in the supplementary materials. while the decoder needs to be specified according to different backbones.
@@ 3.2 Encoder and Decoder
Figure 2: The decoder of ExPert for multi-scale features. The illustration of feature images is visualized by choosing a random slice of the channel dimension. ExPert’s final mask is generated by resizing this mask to the original size.
Transformer backbones can be classified into two types according to whether the scale of features changes. One is the single-scale backbone and the other is the multi-scale backbone. The features of the single-scale backbone keep the same size during the forward propagation while the multi-scale backbone’s features change size. For salient object detection, former research [35, 19, 32] emphasized the importance of multi-scale feature fusion to get finer segmentation masks. Since multi-scale features are crucial to dense prediction for finer details, multi-scale backbones have an advantage over single-scale backbones in segmentation tasks. Therefore we choose the MiT-B4 version of SegFormer [27] as our multi-scale backbone. The encoder of SegFormer has 4 stages of feature extraction and each stage has a different scale of features. The decoder is kept intact as the original one as shown in Fig. 2.
For each stage of feature extraction, denote the output feature of the $i_{th}$ stage as $F^{out}_{i}$. A linear layer $L_{i}$ projects $F^{out}_{i}$ to align the channel dimension for the following concatenation, followed by a bilinear upsampling $\mathcal{U}()$ to align the spatial scale of features. Then all aligned features from different stages are concatenated by $Cat()$ to get $F_{c}$. Finally, two convolution layers, $C_{fuse}$, $C_{pred}$ and the sigmoid function $\mathcal{S}()$ are applied for the mask generation. The mask $M$ is generated as Eq. 2.
@@ 3.3 E-adapter
Figure 3: The detailed structure of E-adapter and E-injector. D-P is the down projection layer, M-L is the median linear layer, U-P is the up projection layer and Dim-P is the dimension projection layer.
To fine-tune the transformer backbone in a parameter-efficient manner, we employ the adapter tuning method, which reduces the dimension of features through a bottleneck design, thereby diminishing the number of trained parameters. As shown in the bottom branch of Fig. 3, we propose the E-adapter, consisting of a down-projection(D-P) layer with parameters $P^{d}_{e}$, a median linear(M-L) layer $L^{m}_{e}$ and an up-projection(U-P) layer with parameters $P^{u}_{e}$. It is noteworthy that the D-P layers of the E-adapter are shared for features of the same spatial scale in order to further diminish the number of trained parameters. The M-L and U-P layers are independent for each E-adapter. The additional low-dimension M-L layers serve to increase the variability of the E-adapter. Define the $i_{th}$ block forward function as $\mathcal{B}_{i}()$, the $(i+1)_{th}$ block’s feature $F_{i+1}$ after the E-adapter is computed as Eq. 4.333The $P\cdot F$ in Eq. 3 represents the linear projection to features F with parameters P.
@@ 3.4 E-injector
Previous works [5, 17] show that side connection modules like adapters can introduce extra information to boost the model’s performance in object detection and segmentation tasks. We design the E-injector to inject external features from other backbones of the same input images as guiding prompts into the encoder for salient object detection. The E-injector’s structure is depicted in the top branch of Fig. 3.
@@
If the number of injected prompt features is $J$, the $j_{th}$ E-injector is composed of a prompt transformation ${Trans}_{j}()$ and a dimension projection (Dim-P) layer with parameters $P^{dim}_{j}$. Since visual prompts might vary in size and shape, a feature transformation ${Trans}_{j}()$ fits the prompt feature $F^{\prime}_{j}$ to the $i_{th}$ layer’s feature $F_{i}$ of the frozen backbone. ${Trans}_{j}()$ is composed of a normalization and a resize operation. The E-injector can receive different transformer features $F^{\prime}_{j}$ including the features from DINO, ViT and BLIP’s vision encoder. The output of the E-injector $F^{inj}_{j}$ is generated as Eq. 5. To better adjust the injector features to the backbone, a learnable scaling vector $\alpha_{j}$ is used to weight them. The $(i+1)_{th}$ block’s feature $F_{i+1}$ is computed as Eq. 6.
$F^{inj}_{j}$ $=P^{dim}_{j}\cdot{Trans}_{j}(F^{\prime}_{j})$ (5)
$F_{i+1}$ $=\mathcal{B}_{i}(F_{i}+F^{ad}_{i}+\sum_{j=1}^{J}F^{inj}_{j}\times\alpha_{j})$ (6)
@@
Finding suitable prompt features is crucial to the quality of salient masks for E-injector. DINO [3] is a self-supervised model trained without labels that exhibits an obvious tendency to focus on objects in an image. Observing the attention maps of DINO reveal coarse masks that are nearly similar to the masks of the salient objects, we assume this kind of object-aware features can aid pre-trained models in locating objects. Besides, considering that our multi-scale backbone has fewer layers of high resolution features compared to single-scale backbones such as ViT, our backbone’s overall perception of an image might be complemented by the features of ViT’s last layer. ExPert uses the features of ViT/B-16 as the auxiliary features for better global perception of images.
@@
While piling up the features might strengthen the visual details, it is still hard to guide the model to recognize the notion of saliency. It is noteworthy that the caption naturally contains the descriptions of the salient objects in an image which are beneficial to SOD. Therefore, we envision that the caption of an image is highly related to the salient objects and contains the salient information. Some Vision Language Models (VLM) like BLIP [13] are trained on large caption datasets. [^4] CLIP [20] is a well-known VLM model trained with millions of image-text pairs. However the text of CLIP is a simple sentence with the class name which is not the caption of the whole image. The resolution of CLIP’s training images is 224*224, the feature is 7*7 with the patch size of 32 which is too small to upsample. Therefore ExPert does not consider CLIP’s features. The features from BLIP’s vision encoder trained with image-text labels, which contain rich semantic information of an image, are injected as semantic-enhanced features by E-injector.
@@
To fully explore the rich semantic information in BLIP, we interact BLIP’s vision features and BLIP’s captions for better focus on salient objects. Although there are no captions or other text information in SOD datasets, we can generate them by the inference results of the BLIP model. For an image $I$, the corresponding caption is generated by BLIP using beam search. The caption is tokenized and embedded by BLIP’s text encoder to get the text embedding $T$. The interacted feature is acquired via the cross-attention between the last layer vision feature $V_{b}$ of BLIP and $T$. In each cross attention layer, $V_{b}$ is projected as the query and $T$ is projected as the key and value by linear projections. Experiments show that one cross-attention layer is enough for the interaction555 More details can be found in the supplementary file. . Since BLIP is trained with image-caption pairs, the alignment of image and text features is not of pixel scale but rather of patch scale. As auxiliary prompt features, too many cross attentions might lead to unexpected noises that harm the final performance.
@@
Since the features from ViT and similar models can refine the mask in detail and the features from BLIP can inject semantic information into ExPert, we combine these two features together for E-injector. The final version of ExPert uses the features from ViT’s last layer and the interacted features by cross-attention from BLIP. Experiments show that the combination prompt version achieves the best performance. Unless specified, ExPert represents our best version.